OpenStudy (sgayton27):

Help me find the sum please! 1 - 1/2 + 1/4 - 1/8 + ... -1/512

OpenStudy (zzr0ck3r):

\(\sum_{n=0}^9(\frac{-1}{2})^n\)

OpenStudy (sgayton27):

yes I keep getting 0.666 (basically 2/3) and it's saying my answer is incorrect.

OpenStudy (anonymous):

(1-1/2)+(1/4-1/8)+(1/16-1/32)... 1/2+1/8+1/32....+1/512 for 6 terms of a geometric progression apply formula

OpenStudy (zzr0ck3r):

right that whole thing works for when we go to infinity, this has finite terms...

OpenStudy (anonymous):

Is it an infinite GP?

OpenStudy (zzr0ck3r):

I just noticed that it didnt have ..... at the end.

OpenStudy (sgayton27):

yeah it ends at -1/512

OpenStudy (sgayton27):

I believe there are 10 terms.

OpenStudy (anonymous):

then first term is 1 and the common difference is -1/2 S = 1* (1-(-1/2)^10)/(1-(-1/2))

OpenStudy (sgayton27):

yes and I get 0.666 but it says my answer is incorrect, maybe the hw is incorrect

OpenStudy (anonymous):

1023/1536

OpenStudy (sgayton27):

wait yeah its 1023/1536!

OpenStudy (sgayton27):

@aviz how did you get to that answer?

OpenStudy (anonymous):

512 = 2^9 so the progression is 1 + (-1/2) + (-1/2)^2... +(-1/2)^9 So there are 10 terms Formula for sum of finite GP = a (1-r^n)/(1-r) a= first term, here 1 r= common ratio = -1/2 n =number of terms 10 S= 1*(1-(-1/2)^10)/(1-(-1/2)) = (1-(1/1024))/(1+1/2) = 1023/1536

OpenStudy (anonymous):

@shamim - I think it is (1-r^n) not (1-r)^n in the numerator

OpenStudy (shamim):

ok

OpenStudy (shamim):

when it will b r^n-1 @aviz